Crew 281 Commander’s Report 15-05-2023
SOL: 1
Commander: Megan Kane
Sol 1, the first day of Team Pegasus’s mission. The crew started waking at dawn, excited for the first day on Mars. Many activities were planned and conducted. Some went as planned, others not so much. However today was a success as we learned, adapted and began settling into the mission.
During our first EVA’s we tested the geotechnical equipment and the procedures that are the bedrock of our crew scientist’s mission. Learning a lot in the process. A soup lunch was followed by an afternoon of habitat tasks and experiment setup culminating in a delicious dinner of quesadillas dipped in the leftover soup.
After a full evening of filing reports and consolidating photos we plan to rest before launching into Sol 2.
Commander Kane signing off.
Journalist Report – February 5th
Crew 274 Journalist Report 2-5-2023
Sol: 7
Author: Tony DiBernardo, Crew Journalist
Title: The Long Hike
Crew 274 woke up this morning at 7:30am to the song “Take Me Home, Country Roads.” by John Denver. The crew got their surveys done, then made sure to get a decent breakfast before their 3-hour EVA.
EVA Activity
Today, HSO Pender and Hab Specialist O’Hara conducted a 3-hour EVA past the supply cache and back. The original EVA plan was to hike an hour to the supply cache, resupply their personal food and water from the cache, then hike 30 more minutes before turning around and hiking 1.5 hours back to the hab. They reached the cache early at the 47-minute mark due to them taking flat roads rather than the originally planned rocky terrain, but eventually found a more consistent and realistic pace on the way back. Supporting their EVA was Commander Guthrie and Journalist DiBernardo. They performed a “leap-frog” maneuver during the EVA, where they would drive the rover past the hiking crew and stop just before losing sight of them. Then, Pender and O’Hara would hike past the parked rover and just before they got out of site, Guthrie and DiBernardo would catch up in the rover and drive past them, only stopping just before losing sight again. This technique was used to reserve rover battery power as well as ensure a safe EVA for the crew on their longest hike of the mission. The EVA was a success, and apart from a couple of cars and hikers driving and passing by, the EVA went as expected.
GreenHab
Following the exciting germination of one LDEF seed yesterday, Greenhab Officer Hines, did his daily collection and observations of germination from the other seeds.. While no additional progress has been noted in other LDEF seeds, multiple sets of seeds that are yet to germinate were removed from the growth tent, placed in wet paper towels, and moved them into the GreenHab to utilize the higher temperatures and humidity in order to potentially germinate leading into the second half of the mission. Additionally, the cress microgreens have continued to make significant progress, with the red cabbage and broccoli making gradual yet minimal continued germination.
Heliophysics
Crew astronomers Pena and Loy’s observations of heliophysics phenomena were significantly impaired by the site’s cloudy weather. Nevertheless, some useful data on solar dark spots and solar prominences were collected and analyzed. These observations will be very useful for the astronomer’s analysis of peak solar cycle 25. So far, 220,000 image points have been collected, and 50,000 images have been processed and analyzed.
Astronomy
Cloudy weather conditions canceled all planned observations of HADS Variable Star V0799 AUR. Crew Astronomers instead continued their analysis of their current image inventory, stacking and analyzing 12 separate 60-second images in the visual filter. So far, HADS Variable Star V0799 AUR is showing signs of dimming, which provides good data to the astronomers’ variable star report.
Stress Mitigation Study
Last night, the crew began their participation in Crew Engineer Lojeks’ study of tracking the stress levels of six of the crew members while practicing breathing and mindfulness techniques as a potential mitigation of stress in analog environments. He spoke with the crew about why he is pursuing this study, how to enter into this type of meditation and guided them through a 15-minute breathing exercise.
Commander Report – December 8th
Sol: 11
Summary Title: EVA #15 Balloon Launch
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:
Crew 269 enjoyed another day of beautiful weather on Mars with sunny skies, very calm winds, and mild temperatures. The crew had a quick breakfast of cereal, dried fruit, and homemade sourdough bread, and HSO collected health and safety stats. The crew held a morning meeting to assess effects to the day’s EVA plan due to the melt from yesterday’s snow continuing to create muddy conditions around the Hab. After consult with Tier 1 mission support, the crew’s plan for a single afternoon EVA to launch the weather balloon payload was confirmed and the crew prepared for the experiment. CDR, XO, and Green Hab Officer embarked on EVA #15 and immediately began to set up and fill the weather balloon with helium in the area in front of the Hab. Once the balloon fill was complete, the crew walked the balloon around to the north side of the Hab just beyond the observatory to conduct the launch. A technical hiccup with the telemetry was successfully resolved, and the balloon launched at 1:45 PM. The crew immediately began receiving telemetry from the weather balloon and was in contact with the balloon over eastern Utah and western Colorado for several hours. During the EVA, Engineer received further instruction from the PI for the mirror experiment who wished to have one more microscope analysis before packing up the mirrors tomorrow for return to Aerospace. IVA crew requested and received permission from Tier 1 mission support to add an additional objective to EVA #15 to retrieve the mirror samples from their exposure location across Cow Dung Rd. from Marble Ritual and return them to the Hab for analysis. After the EVA, Engineer analyzed the mirror samples under the Aerospace microscope in the Science Dome and Green Hab Officer tended the plants in the Green Hab. For dinner, the crew enjoyed shepherd’s pie. The crew plans to break simulation tomorrow morning at 5:30 AM to complete a live STEM event, photography/videography session, and end of mission cleanup activities.
Commander Report – December 7th
Sol: 10
Summary Title: Snow Day and EVA #14
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:
Crew 269 awoke to a winter wonderland landscape this morning and snowy conditions. The crew enjoyed a quiche bake for breakfast and conducted a morning meeting to discuss how the weather would affect the day’s EVA plans. After a consultation with Tier 1 mission support, the crew decided to shorten the morning EVA #14 and continue to send 2 crew to perform repair operations on the mirror coating experiment payload. HSO and Engineer traversed on foot through the muddy terrain to repair the mirror payload and return swiftly to the Hab, while Green Hab Officer lead IVA comms. During the EVA, CDR and XO discussed end of mission reporting requirements and began to create content for those reports while Technology Officer continued to work data processing on the Aerospace computers. In the afternoon, HSO collected health and safety stats, crew completed exercise protocols, and Green Hab Officer tended to the plants in the Green Hab. Power was again limited due to weather conditions and the crew followed Tier 1 support instructions to conserve power. For dinner, the crew enjoyed black bean burgers on homemade sourdough buns with microgreens from the Green Hab.
Commander Report – December 6th
Sol: 9
Summary Title: EVAs #12 and #13
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:
Crew 269 experienced another day of great weather on Mars with mostly mild temperatures, sunny skies, and light high-altitude clouds, with morning to midday cloud cover and a sprinkle of rain. The crew awoke early this morning and enjoyed a quick breakfast of cereal, oatmeal, and dried fruit. After a morning planning meeting, XO, HSO, and Technology Officer conducted the morning EVA (#12) for the purpose of flying the drone and to continue demonstrating the VR/AR headset in the area around the Hab and along Cow Dung Rd. to Marble Ritual. During the morning EVA, Engineer and Green Hab Officer exchanged the furnace filter and changed the battery on space suit #6. Upon return of the EVA #12 crew, and after a quick lunch, CDR, Engineer, and Green Hab Officer embarked on EVA #13 south on Cow Dung Rd. and then on foot on Rte 1101 to Barrainca Butte for the purpose of regolith sample collection, radiation environment measurements, and handheld ham radio comm checks. Throughout the day, the crew was updated on the status of the power system by Tier 1 support and entered various levels of power conservation as directed. In the afternoon, HSO collected health stats, crew completed exercise protocols, and conducted a test of the weather balloon payload scheduled to launch later this week. For dinner the crew enjoyed pesto pasta with vegetables with homemade sourdough bread.
Commander Report – December 5th
Sol: 8
Summary Title: EVAs #10 and #11
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:
Crew 269 experienced another day of beautiful weather on Mars with mild temperatures, sunny skies, and light high-altitude clouds. Green Hab Officer made blueberry pancakes for breakfast with bourbon maple syrup. After a morning planning meeting, CDR, XO, and HSO conducted EVA #10. The first objective of the EVA was to traverse on foot to retrieve the mirror coating experiment samples from their exposure location across Cow Dung Rd. from Marble Ritual and return the samples to the airlock for analysis under microscope in the Science Dome by Engineer. The EVA then proceeded on rovers to Kissing Camel Ridge, then on foot up to Hab Ridge Rd. for the purpose of regolith sample collection and radiation environment monitoring. After EVA #10 crew returned to the Hab, Engineer and Green Hab Officer embarked on EVA #11, traversing on foot to return the mirror coating experiment samples to their exposure location across Cow Dung Rd. from Marble Ritual. The EVA then proceeded on foot to Sagan St. for the purpose of a handheld ham radio range test with the Hab. The handheld ham radios maintained clear communications on the 2 meter band at a further distance from the Hab than was possible using the EVA radios. During the afternoon EVA, HSO conducted in-depth troubleshooting of the Hab toilet (details can be found in the operations report for today), and Technology Officer conducted troubleshooting of the VR/AR headset in preparation for future EVAs. In the early evening, Green Hab Officer harvested microgreens, HSO collected crew health stats, and the crew took turns completing their exercise protocols. For dinner, the crew enjoyed chicken jambalaya with homemade sourdough bread and a microgreen salad. After dinner, the crew met to discuss prioritizing the remaining mission objectives to ensure maximum science return with the time remaining.
Commander Report – December 4th
Sol: 7
Summary Title: EVA #9 and Suit Troubleshooting
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:
Crew 269 experienced another day of beautiful weather on Mars with mostly mild temperatures, sunny skies, and light high-altitude clouds. The crew enjoyed an egg, cheese, and vegetable casserole for breakfast. After a morning planning meeting, HSO collected crew health stats. CDR, Technology Officer, and Green Hab Officer conducted a short EVA in the early afternoon to deploy the ham radio antenna and fly two drone flights around the Hab and around the mirror coating experiment sample location. Fortunately, winds were light today, so the drone was able to successfully complete both flights. During the EVA, Engineer and HSO completed troubleshooting of the two-piece EVA suits and determined that the power strip the charging cables were plugged into had gone bad. They completed further tests with various combinations of suits, batteries, and chargers to ensure all six suits are now charging and holding charge properly. The crew relaxed a bit in the afternoon, reading, knitting, working on art projects, and sharing photos and stories. For dinner, the crew enjoyed tacos with homemade tortillas and baked a cake to celebrate the midpoint of their mission.
Commander Report – December 3rd
Sol: 6
Summary Title: EVAs #7 and #8
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:
Crew 269 experienced another day of beautiful weather on Mars with mostly mild temperatures, sunny skies, and light high-altitude clouds. The crew enjoyed a breakfast of homemade sourdough bread, cereal, oatmeal, dried fruit, and espresso. After a morning planning meeting, HSO collected crew health stats. CDR, XO, and Engineer conducted the morning EVA #7 to the Special Region and used an EVA tool suite (rake, scoop, special sample bags) to collect regolith samples, being careful to adhere to the restrictions at the site to not disturb any fossils. CDR also collected GPS-tagged radiation dose rate measurements using a handheld dosimeter. The crew noted that this region would be an excellent site to fly the drone on a future Sol. During their return from the Special Region, EVA #7 crew marked a second site for a sample location on Cow Dung Rd. for the EVA #8 crew to return in the afternoon. After a quick lunch, HSO and Green Hab Officer conducted EVA #8 to the site on Cow Dung Rd. marked by the EVA #7 crew, where they collected further regolith samples, and HSO collected further radiation dose measurements with the handheld dosimeter. During the EVA, Technology Officer sent data to Aerospace mission support and received feedback on previously uploaded data. In the afternoon, CDR and Engineer inspected the 2-piece EVA suits. Further information can be found in today’s operations report. Before dinner, the crew completed their montage painting and assembled their inputs for the mid-mission research report. For dinner, the crew enjoyed chicken pot pie.
Commander Report – December 2nd
Sol: 5
Summary Title: EVAs #5 and #6
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:
Crew 269 experienced another day of beautiful weather on Mars with mostly mild temperatures, sunny skies, and light high-altitude clouds. Winds were moderate most of the day, with higher gusts, with a brief drizzle in the late morning. The crew enjoyed a breakfast of cereal, oatmeal, homemade sourdough bread, and espresso. After a morning planning meeting, HSO collected crew health stats. XO and Engineer conducted a short morning EVA (#5) to retrieve the mirror samples to analyze in the Science Dome under a microscope. Technology Officer employed help from the two licensed private/drone pilots on the crew (CDR, HSO) to complete drone flight planning preparations for the afternoon EVA. After a light lunch, CDR, Technology Officer, and HSO embarked on EVA #6, while the Green Hab Officer led IVA Comms. The EVA plan included drone operations with flight plans near the Hab, near the mirror sample deployment location, and at Marble Ritual. The crew successfully completed activation and checkout of the drone near the Hab but decided to call off Flight Plan #1 due to high wind gusts. The crew proceeded to perform their backup EVA plan operations demonstrating the VR/AR headset in the area around the Hab and at the intersection of Hab Rd and Cow Dung Rd. At the mirror deployment location and Marble Ritual, the crew again called off the drone Flight Plans #2 and #3 due to high wind gusts and performed further demonstration of the VR/AR headset in those areas. The crew then returned early to the Hab area; by that time the winds had died down slightly allowing for the crew to proceed with Flight Plan #1 around the Hab after requesting permission from Tier 1 support to modify the original EVA plan. The drone flight was successful, collecting imagery of the Hab area. Finally, the EVA crew retrieved the mirror samples from the airlock and redeployed them at the location across Cow Dung Rd. from Marble Ritual before returning to the Hab. During the EVA, Engineer analyzed the mirror samples under the Aerospace microscope in the Science Dome. In the afternoon, Green Hab Officer harvested carrot sprouts and the crew continued their montage painting. The crew ended the day with a pizza party.
Commander Report – December 1st
Sol: 4
Summary Title: Space Suit Repair & EVA #4
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:
Crew 269 experienced another day of beautiful weather on Mars with sunny skies, light high-altitude clouds and calm winds throughout the morning and into the afternoon. In the late afternoon, winds picked up and clouds began to move in. The crew enjoyed a breakfast of cereal, oatmeal, homemade sourdough bread, and espresso. After a morning planning meeting, HSO collected crew health stats. Technology Officer employed help from the two licensed private/drone pilots on the crew (CDR, HSO) to complete drone flight planning preparations for the afternoon EVA. After exercise and a light lunch, CDR, Technology Officer, and HSO embarked on EVA #4, while the XO led IVA Comms. The EVA plan included drone operations along Cow Dung Rd to Galileo Rd, but early technical challenges with the drone caused the crew to deviate from the scheduled EVA plan, remaining close to the Hab to troubleshoot. Due to this deviation, the crew was called back early to the Hab. During the EVA, Engineer and Green Hab Officer successfully completed repair operations on EVA suit 7 in the RAM. All crew members enjoyed assisting the Green Hab Officer in planting carrots and onions in the Green Hab in the afternoon. Before dinner, CDR, Technology Officer, and HSO replanned the drone flights for tomorrow’s requested morning EVA. The three ham radio operators on the crew (CDR, Green Hab Officer, and HSO) studied how to deploy the ham radio antenna for tomorrow’s requested afternoon EVA. The crew enjoyed a dinner of mac and cheese with homemade sourdough bread and peanut butter chocolate chip cookies.